# Network Formation and Dynamics Among Multi-LLMs

### v3

@@ Results
We generated networks of various sizes, ranging from $n=10$ to $n=100$, to explore the relationship between the network size ($n$) and two key metrics: the average shortest path length ($L$) and the average clustering coefficient ($C$). For this analysis, we considered values of $\beta$ set at 0.25, 0.5, and 0.75, with a fixed $k=5$ to serve as a consistent parameter. Visual representations and average clustering coefficients are presented in Figures 4(a), 4(b) and 4(c).
However, when directly compared with the Watts-Strogatz model, the networks generated by the LLMs do not precisely replicate the characteristics of Watts-Strogatz networks for the corresponding the rewiring probabilities ($\beta$). As illustrated in right panels of Figures 4(a), 4(b) and 4(c), we find weak evidence that the LLM-generated networks share the same average shortest path length as the Watts-Strogatz model for the rewiring probabilities ($\beta$) of 0.25, 0.5, and 0.75, with $P<0.1$ from a t-test comparing the average shortest path lengths. Additionally, LLM-generated networks exhibit a larger average clustering coefficient than those of the Watts-Strogatz model for the same rewiring probabilities $\beta$, also with a significance level of $P<0.1$ in the t-test comparisons. This discrepancy suggests the differences in the network structure and connectivity patterns between the LLM-generated networks and the classical Watts-Strogatz model.
@@
To quantify how LLM-generated networks resemble Watts-Strogatz networks, we fit the estimated $\hat{\beta}$ values for each LLM-generated network. [^10] We conducted a binary search to identify the $\hat{\beta}$ values for which the Watts-Strogatz networks’ average clustering coefficients match those of the LLM-generated networks. In Figures 5(d), 5(e) and 5(f), we plot the estimated values for $\hat{\beta}$ for each value of $\beta$ and each temperature. Here $P$-values result from a t-test comparing with the average shortest path length of Watts-Strogatz with rewiring probability $\hat{\beta}$. These results show that while the average shortest path lengths are not identical, they are sufficiently close, with the differences not being statistically significant at the $0.1$ level for most temperature settings.
@@
Decisions on Real-World Networks
@@
To investigate the behavior of LLMs in real-world network formation contexts, we employ the Facebook100 dataset (Traud et al., 2012) and a discrete choice modeling framework (Overgoor et al., 2019; McFadden, 1972). This dataset comprises “friendship” networks from one hundred American colleges and universities, captured at a specific moment in time from Facebook’s online social network. We model the network formation process as a discrete choice process, wherein nodes are sequentially prompted to form connections from a set of available alternatives (see Methods and Materials).
We use datasets that refer to three US colleges – Caltech36, Swarthmore42, and UChicago30. They represent different sizes of networks: $n=769$, $n=1,659$, and $n=6,591$ nodes, respectively. For the Caltech36 ($n=769$) and Swarthmore42 ($n=1,659$) datasets, we consider all of the network nodes to be nodes that have to make a choice. For the UChicago30 dataset, we consider a randomly sampled subset of 2,000 nodes because of the limited context window of GPT models. We set the size of the alternative sets $A_{t}$ to 20, and thus, the accuracy of random guessing is 5%.
Figure 6. Regression results for real-world network experiments. Multinomial logit coefficients for three networks from the Facebook100 dataset. The standard errors of the estimates are shown with error bars. The full regression tables can be found in SI Section 1. All coefficients are statistically significant, where the null hypothesis corresponds to the respective parameter being equal to 0.
We regress network formation decisions on standardized scores reflecting the three micro-level principles. We present the regression results in Figure 6.111111More detailed results can be found in Table S1 in SI Section 1. Our analysis reveals that the degree, the number of common neighbors, and the number of shared interests are all statistically significant factors influencing the choices of LLMs ($P<0.001$) across all datasets. Furthermore, we find that LLMs place the greatest emphasis on the number of common attributes ($\hat{\theta}_{\mathrm{H}}$), followed by the number of common neighbors ($\hat{\theta}_{\mathrm{TC}}$), and then the degree ($\hat{\theta}_{\mathrm{PA}}$).
These results suggest that LLMs exhibit a strong preference for homophily, followed by a tendency towards triadic closure, and finally, preferential attachment. Specifically, the relationship $\hat{\theta}_{\mathrm{H}}>\hat{\theta}_{\mathrm{TC}}>\hat{\theta}_{\mathrm{PA}}$ holds across all examined datasets and temperature settings.
Further insights are provided in SI Section 4, where we provide the reasoning analysis. The analysis points to a marked emphasis on homophily, indicating that nodes with similar characteristics—such as student/faculty status, high school, and graduation year—are more likely to connect. In contrast, triadic closure and preferential attachment have a relatively lesser impact.
Our study also includes a comparison of LLMs’ predictive capabilities against random guessing in a link prediction framework. We assess the percentage increase in accuracy, defining accuracy as the probability of selecting an existing neighbor from a set of alternatives, as opposed to a random selection from 20 possible options. As detailed in SI Section 1, LLMs significantly outperform random guessing in accuracy.
Furthermore, we investigate whether LLMs reinforce community structure post-network formation. We run the Louvain algorithm 10 times for both the initial graph $G_{1}$ and the final graph $G_{T}$ and perform a t-test between the resulting modularity scores (see SI Section 1). We observe that the modularity increases in all cases and is statistically significant. This suggests that LLMs make choices that naturally yield a reinforcement of the community structure.
Additionally, to explore the “small-world” phenomenon, we measure the percent change in average shortest path length between the pre- and post-decision networks (detailed in SI Section 1). Our findings reveal a minor decrease in average path length in most cases, suggesting a slight strengthening of the small-world phenomenon.
@@
In summary, our analysis demonstrates that while homophily, triadic closure, and preferential attachment are integral to the network formation behaviors of LLMs, homophily is the dominant factor. Furthermore, both community structure and the small-world phenomenon are reinforced after network formation decisions of LLMs.
@@ Discussion
Our study conducted a comprehensive evaluation of LLMs’ network formation preferences, encompassing both micro-level network principles such as preferential attachment, triadic closure, and homophily and macro-level network principles like community structure and the small-world phenomenon. Our findings indicate that networks generated by multiple LLMs do exhibit these properties, particularly when primed with network statistics such as the number of mutual friends or the degree of alternatives. Furthermore, through discrete choice modeling, we explored the emergence of these properties in real-world networks. Our real-world network results reveal that the agents’ selections are predominantly driven by homophily, followed by triadic closure, and then preferential attachment. Additionally, we observed a strengthening of community structure in real-world network experiments.
Our study enhances our understanding of the behavior of multi-LLMs. For example, our findings reveal different strengths of network formation properties, suggesting that when LLMs are employed to coordinate social networks in social or work environments, they may exhibit behaviors similar to humans. This insight contributes to developing socially aware LLMs that align with human preferences, with potential applications ranging from chatbots and virtual assistants to collaborative systems and social media platforms.
Furthermore, our study may pave the way for innovative applications of LLMs in network science and computational social science research. Given that LLMs adhere to fundamental network principles and exhibit behaviors akin to human network formation, they hold significant potential for use in simulations of social systems and agent-based modeling. As LLM agents gain increasing ability to exhibit more complex and nuanced behaviors, such as negotiation, collaboration, and competition, this presents opportunities to replicate networked environments like marketplaces or organizational structures. In these settings, LLMs could effectively model decision-making processes and interaction dynamics.
Looking ahead, we propose three promising directions for future research. First, we propose investigating LLM behavior in more complex interactions, such as simulated dialogs or specific environments like companies, schools, or parties. By examining LLMs in these contexts, we can better understand their network formation preferences and how they adapt to different social dynamics. This research could shed light on the nuances of LLM behavior in realistic settings and inform their integration into diverse social environments. Second, we suggest exploring the application of our findings in specific settings, particularly in organizational contexts. For example, considering the potential use of LLMs to assist HR professionals, our study’s insights could guide the alignment of LLMs with organizational goals. By leveraging network information, LLMs could help HR departments identify their firms’ best fit or talent, optimizing recruitment and talent management processes. Finally, we propose using our methods to create realistic synthetic networks. These synthetic datasets can serve as benchmarks for evaluating graph learning methods, thus addressing the scarcity of existing graph benchmarks. By adjusting parameters like temperature, we can generate diverse networks to test graph neural network performance under various conditions. In addition, our approach can generate artificial data resembling real-world data while adhering to privacy regulations. This has the potential to expedite research in graph learning and trustworthy machine learning by providing realistic network datasets generated by LLMs.
1. Methods
@@
ALGORITHM 1 Prompt used to implement $\mathcal{Q}(A_{t},i_{t},\delta)$.
@@
We employ multiple temperatures to account for the variability in response generation by LLM systems, which is also observed in classical statistical models of network formation (Jackson, 2008). Our study conducts experiments using three temperatures: 0.5, 1.0, and 1.5.
@@
Summarizing the Reasoning of the LLM Agents
In addition to analyzing the innate preferences of LLM agents, we employ an interview-like approach to probe the rationale behind their decisions. Following the LLMs’ decisions on network formation, we inquire about their reasoning for making those choices. We then aggregate all the provided rationales using GPT models (GPT-3.5 for the synthetic networks and GPT-4 for the real-world networks) by asking it to identify the common themes, resulting in the creation of multiple categories. To determine the category to which each LLM’s reasoning belongs, we subsequently ask the GPT models to classify each rationale into the appropriate category.
@@
The algorithm of generating LLM networks is described as follows:
@@
where $i_{1},\dots,i_{T}$ are the chooser nodes (i.e., the LLM agents who want to form a link), and $j_{1},\dots,j_{T}$ are the nodes which are chosen from the alternative sets $A_{1},\dots,A_{T}$. The likelihood function is convex, and we optimize it with the L-BFGS-B method (Liu and Nocedal, 1989). The standard errors of the coefficients are approximated as $\sqrt{-H^{-1}}$ where $H$ is the Hessian matrix of the log-likelihood at $(\hat{\theta}_{\mathrm{PA}},\hat{\theta}_{\mathrm{TC}},\hat{\theta}_{\mathrm{H}})$ (cf. (Overgoor et al., 2019; Train, 2009)).
@@
Data and code are available at GitHub: https://github.com/papachristoumarios/llm-network-formation.

### v4

@@ Results
We generated networks of various sizes, ranging from $n=10$ to $n=100$, to explore the relationship between the network size ($n$) and two key metrics: the average shortest path length ($L$) and the average clustering coefficient ($C$). For this analysis, we considered values of $\beta$ set at 0.25, 0.5, and 0.75, with a fixed $k=5$ to serve as a consistent parameter. Visual representations and average clustering coefficients are presented in Figures 5(a), 5(b) and 5(c).
However, when directly compared with the Watts-Strogatz model, the networks generated by the LLMs do not precisely replicate the characteristics of Watts-Strogatz networks for the corresponding the rewiring probabilities ($\beta$). As illustrated in right panels of Figures 5(a), 5(b) and 5(c), we find weak evidence that the LLM-generated networks share the same average shortest path length as the Watts-Strogatz model for the rewiring probabilities ($\beta$) of 0.25, 0.5, and 0.75, with $P<0.1$ from a t-test comparing the average shortest path lengths. Additionally, LLM-generated networks exhibit a larger average clustering coefficient than those of the Watts-Strogatz model for the same rewiring probabilities $\beta$, also with a significance level of $P<0.1$ in the t-test comparisons. This discrepancy suggests the differences in the network structure and connectivity patterns between the LLM-generated networks and the classical Watts-Strogatz model.
@@
To quantify how LLM-generated networks resemble Watts-Strogatz networks, we fit the estimated $\hat{\beta}$ values for each LLM-generated network. [^8] We conducted a binary search to identify the $\hat{\beta}$ values for which the Watts-Strogatz networks’ average clustering coefficients match those of the LLM-generated networks. In Figures 6(d), 6(e) and 6(f), we plot the estimated values for $\hat{\beta}$ for each value of $\beta$ and each temperature. Here $P$-values result from a t-test comparing with the average shortest path length of Watts-Strogatz with rewiring probability $\hat{\beta}$. These results show that while the average shortest path lengths are not identical, they are sufficiently close, with the differences not being statistically significant at the $0.1$ level for most temperature settings. Finally, as Figure 6(g) shows, the relation $L\sim\log(n)$ holds for different LLM models and environments.
@@
Decisions on Real-World Networks with Heterogeneous Agents
We investigate the behavior of LLMs in real-world network formation contexts with four datasets in two differing real-world domains. Despite the significant advancements in social network analysis over recent years, the availability of fully complete and comprehensive network datasets remains exceptionally rare (Yuan et al., 2018). We employ three datasets from the Facebook100 collection (Traud et al., 2012) and the telecommunication (Andorra) and the employment (MobileD) datasets from (Yuan et al., 2018). The Facebook100 data correspond to “friendship” networks from one hundred American colleges and universities, captured at a specific moment from Facebook’s online social network. The Andorra dataset contains nationwide call records in Andorra from July 2015 to June 2016, where calls correspond to mutual calls between Andorran residents containing information about the caller’s and the callee’s location, phone type, and usage. Finally, the MobileD dataset corresponds to a company network where relations correspond to call or text communication, and each employee is either a manager or a subordinate.
For all network datasets, the agents have heterogeneous profiles (i.e., profiles with different features) whose statistics (degree distribution, clustering coefficient distribution, assortativity) we report in Figure 7.
@@
To infer the models’ tendencies, we employ a discrete choice modeling framework (Overgoor et al., 2019; McFadden, 1972). Specifically, we model the network formation process as a discrete choice process, wherein nodes are sequentially prompted to form connections from a set of available alternatives (see Methods and Materials).
Number of Samples and Alternatives
For the three datasets from Facebook100 are Caltech36 ($n=769$) Swarthmore42 ($n=1,659$), and UChicago30 ($n=6,591$), we set the number of alternatives to be $A=15$ and randomly sampled from the existing network. For the UChicago30 dataset, we consider a randomly sampled subset of $N=2,000$ nodes because of the limited context window of the LLM models. For Andorra ($n=32,812$) and MobileD ($n=1,982$), we set the number of alternatives to $A=5$ and consider a randomly sampled subset of $N=1,000$ nodes.
Figure 7. Distributions of real-world datasets analyzed in our study, including degree, clustering coefficients, and the assortativities of the attributes included in the datasets.
Model Preferential Attachment ($\hat{\theta}_{\mathrm{PA}}$) Homophily ($\hat{\theta}_{\mathrm{H}}$) Triadic Closure ($\hat{\theta}_{\mathrm{TC}}$) Log Likelihood AIC
Caltech36 ($n=769$ nodes, $m=33,312$ edges, $N=769$ samples, $A=15$ alternatives each)
GPT-3.5 0.20*** (0.002) 0.65*** (0.005) -0.06 (0.006) -2,088.21 4,184.41
GPT-4o Mini 0.34*** (0.006) 2.13*** (0.03) 0.44*** (0.02) -1,201.27 2,410.55
GPT-4 (gpt-4-1106-preview) 0.41*** (0.01) 1.95*** (0.02) 0.59*** (0.01) -1,377.47 2,762.94
Claude 3.5 Sonnet 0.46*** (0.005) 0.55*** (0.01) 0.55*** (0.007) -1,748.19 3,504.38
Llama 3 70b Instruct 0.28*** (0.006) 2.43*** (0.02) 0.84*** (0.01) -809.57 1,627.15
Swarthmore42 ($n=1,659$ nodes, $m=122,100$ edges, $N=1,659$ samples, $A=15$ alternatives each)
GPT-3.5 0.19*** (0.008) 0.47*** (0.01) 0.00 (0.009) -4,484.45 8,976.90
GPT-4o Mini 0.27*** (0.21) 2.22*** (0.78) 0.57*** (0.43) -1,899.09 3,806.19
GPT-4 (gpt-4-1106-preview) 0.18*** (0.003) 1.62*** (0.006) 0.65*** (0.002) -2,838.33 5,684.66
Claude 3.5 Sonnet 0.36*** (0.002) 0.75*** (0.006) 0.55*** (0.004) -3,563.02 7,134.03
Llama 3 70b Instruct 0.39*** (0.003) 2.31*** (0.005) 0.62*** (0.004) -1,820.26 3,648.52
UChicago30 ($n=6,951$ nodes, $m=416,206$ edges, $N=2,000$ samples, $A=15$ alternatives each)
GPT-3.5 0.22*** (0.001) 0.48*** (0.004) -0.02 (0.0005) -8,157.38 16,322.77
GPT-4o Mini 0.27*** (0.005) 2.22*** (0.019) 0.57*** (0.011) -1,899.09 3,806.19
GPT-4 (gpt-4-1106-preview) 0.23*** (0.001) 2.00*** (0.005) 0.41*** (0.002) -3,444.33 6,896.67
Claude 3.5 Sonnet 0.43*** (0.003) 0.78*** (0.005) 0.39*** (0.002) -6,604.77 13,217.54
Llama 3 70b Instruct 0.43*** (0.007) 2.57*** (0.014) 0.32*** (0.005) -3,689.00 7,386.00
Andorra ($n=32,812$ nodes, $m=513,931$ edges, $N=1,000$ samples, $A=5$ alternatives each)
GPT-3.5 0.53*** (0.001) 0.21* (0.01) -0.24*** (0.002) -1,712.91 3,433.83
GPT-4o Mini 0.54*** (0.004) 3.47*** (0.06) -0.09* (0.01) -1,002.11 2,012.22
GPT-4 (gpt-4-1106-preview) 0.21 (0.20) 3.45*** (1.16) -0.22 (0.25) -1271.20 2622.30
Claude 3.5 Sonnet 0.54*** (0.003) 1.94*** (0.009) -0.15*** (0.003) -1,541.77 3,091.55
Llama 3 70b Instruct 0.38*** (0.003) 3.92*** (0.02) -0.04 (0.01) -985.95 1,979.91
MobileD ($n=1,982$ nodes, $m=25,470$ edges, $N=1,000$ samples, $A=5$ alternatives each)
GPT-3.5 1.06*** (0.003) -0.94*** (0.009) -0.02 (0.001) -1,663.42 3,334.84
GPT-4o Mini 1.38*** (0.02) -0.85*** (0.02) 0.87*** (0.01) -880.39 1,768.78
GPT-4 (gpt-4-1106-preview) 0.42*** (0.01) -1.84*** (0.009) 0.94*** (0.003) -1,321.01 2,650.02
Claude 3.5 Sonnet 0.71*** (0.009) -2.44*** (0.02) 1.13*** (0.005) -1,197.92 2,403.83
Llama 3 70b Instruct 1.04*** (0.005) -0.36** (0.01) 0.71*** (0.002) -1,269.42 2,546.83
Note: *: $P<0.05$, **: $P<0.01$, ***: $P<0.001$
Table 1. Effect sizes for real-world networks from Facebook100 (Traud et al., 2012) and Andorra dataset (Yuan et al., 2018) for several LLMs for temperature set to 0.5. An earlier version of this paper included experiments with gpt-4-1106-preview beyond gpt-4o-mini. We have also added these experiments for completeness.
Regression Coefficients
We regress network formation decisions on standardized scores reflecting the three micro-level principles. We present the regression results in Table 1.999More detailed results can be found in Table S.2 in Appendix A. First, we observe a dominant effect of homophily across all datasets and models. The coefficients for homophily ($\hat{\theta}_{\mathrm{H}}$) are consistently the largest and highly significant ($P<0.05$) in almost all cases. For instance, in the Caltech36 dataset, the homophily coefficients for GPT-3.5, GPT-4, and Llama 3 70b Instruct are 0.65, 1.95, and 2.43, respectively ($P<0.001$). The emphasis on homophily suggests that LLMs, much like humans, prioritize forming connections based on shared characteristics.
Second, we find that preferential attachment plays a secondary role in the network formation decisions of LLMs. While the coefficients for preferential attachment ($\hat{\theta}_{\mathrm{PA}}$) are generally positive and statistically significant across most models and datasets, they are notably smaller than those for homophily. For example, in the Swarthmore42 dataset, GPT-3.5 and Llama 3 70b Instruct have preferential attachment coefficients of 0.19 and 0.39, respectively ($P<0.001$). This suggests that while LLMs do consider the degree of potential connection nodes—favoring connections to well-connected nodes—the influence of this factor is less evident compared to homophily.
Finally, the influence of triadic closure appears to vary across different datasets and models. In most cases, the coefficients for triadic closure ($\hat{\theta}_{\mathrm{TC}}$) are positive and significant, indicating that LLMs consider the number of mutual connections when forming new links. However, in some instances, such as with GPT-3.5 on the Andorra dataset, the triadic closure coefficient is negative ($-0.24$) and significant, suggesting a structure-dependent role of this principle as shown by the low clustering coefficient of the network, which is dominated by preferential attachment, as also shown in Figure 7. This variability implies that while triadic closure is a factor in LLMs’ decision-making, its impact may be influenced by the specific characteristics of the dataset or the model used.
@@
In summary, our analysis demonstrates that while homophily, triadic closure, and preferential attachment are integral to the network formation behaviors of LLMs, homophily is the dominant factor.
@@ Discussion
In this study, we conducted a comprehensive evaluation of LLMs’ network formation preferences, examining both micro-level network principles—such as preferential attachment, triadic closure, and homophily—and macro-level network properties like community structure and the small-world phenomenon. Our findings indicate that networks generated by multiple LLMs exhibit these properties, particularly when the models are primed with network statistics like the number of mutual friends or the degrees of potential connections. Furthermore, using discrete choice modeling, we explored the emergence of these properties in simulations based on real-world networks. Our results reveal that the LLM agents’ selections are predominantly driven by homophily, followed by triadic closure and preferential attachment.
On the one hand, our study enhances our understanding of how multiple LLMs behave in networked settings. Specifically, our findings reveal varying strengths in network formation properties among LLMs, suggesting that when these models are employed to coordinate social networks in social or work environments, they may exhibit human-like behaviors. This has important implications for applications like agent-based modeling, where realistic simulation of human behavior is crucial. Traditionally, agent-based models rely on simplified rules or heuristics to represent individual behaviors, which may not capture the complexity of human decision-making. By incorporating LLMs as agents, we can simulate more nuanced and context-aware interactions that closely resemble human social behavior, without the need to rigidly specify decision rules or heuristics.
Limitations and Boundary Conditions
On the other hand, our results suggest that we should exercise caution when leveraging LLMs in networking scenarios. The model family, configuration, and prompts can subtly affect the models’ behavior, resulting in qualitatively similar but quantitatively different outcomes.
For example, newer models such as GPT-4 and Claude 3.5 exhibit stronger biases compared to prior models such as GPT-3.5. For instance, in the preferential attachment principle, newer models such as GPT-4 and Claude 3.5 have stronger biases – i.e., connect to highest-degree-nodes yielding star-like networks – compared to GPT-3.5 and LLama 3, which had weaker biases – i.e., connect to high-degree-nodes. Similar results can be found in homophily, as larger biases towards homophily, and triadic closure. Thus, even though LLMs exhibit these principles, we should be cautious of such biases when designing simulations.
Similarly, in our experiments with real-world data, we find an interesting phase transition from homophily to heterophily: In the Facebook100 data, the LLMs generally exhibited positive biases toward homophily ($\hat{\theta}_{H}>0$). However, in employment networks, agents were either managers or subordinates, and we discovered that LLM agents were heterophilous in such a case ($\hat{\theta}_{H}<0$), which aligns with career advancement dynamics (i.e. employees want to form links with managers because of better career prospects).
The above underscores the need for researchers to provide oversight and ensure that LLM behaviors align with human expectations when employing them in scientific research methods, such as agent-based modeling and even prototypical human subject research with LLMs.
Thus, although we find that LLMs resemble human network formation behaviors, we should consider whether these models should exhibit such behaviors when serving as assistants to humans in work and social lives. Biases like homophily, triadic closure, or preferential attachment may lead to network structures that overemphasize certain individuals or fragment information flow. Humans form networks with communities and central nodes partly due to limited social capacities; however, LLMs do not share these limitations.
Therefore, when used as social assistants, LLMs may not necessarily need to mirror human networking behaviors and could be personalized to promote more equitable and efficient information dissemination. Our study highlights the need for more deliberate efforts to align LLM behavior in this domain.
Future Research Directions
Looking ahead, we identify three promising directions for future research: First, investigating LLM behavior in more complex interactions, such as simulated dialogues, could provide deeper insights. By examining the specific dialogues of LLMs during interactions, we can better understand their network formation preferences and how they adapt to different social dynamics. Second, we could explore how LLMs can be integrated into real-world settings, such as social media platforms. LLM-assisted bots might be employed to facilitate interactions, break echo chambers, and moderate democratic discussions (see, for example, (Papachristou et al., 2023)).
Finally, we can use our methods to create realistic synthetic networks. These synthetic datasets can serve as benchmarks for evaluating graph learning methods, thereby addressing the scarcity of existing graph benchmarks (cf. (Palowitch et al., 2022)). By adjusting parameters like temperature and environmental settings or selecting different models, we can generate diverse networks to test graph neural network performance under various conditions. Importantly, our approach allows for the generation of artificial data that resembles real-world data while adhering to privacy regulations.
@@
# Task You are located in a school. Your task is to select a set of people to be friends with. # Profile Your profile is given below after chevrons: <PROFILE> { "name" : "Person 0", "favorite subject" : "Chemistry", "neighbors" : ["Person 3", "Person 432", "Person 4", "Person 3", "Person 32"] } </PROFILE> # Candidate Profiles The candidate profiles to be friends with are given below after chevrons: <PROFILES> [ { "name" : "Person 1", "favorite subject" : "Mathematics", "neighbors" : ["Person 3", "Person 4", "Person 23", "Person 65"] }, { "name" : "Person 33", "favorite subject" : "History", "neighbors" : ["Person 342", "Person 2", "Person 12"] }, ... ] </PROFILES> # Output The output should be given a list of JSON objects with the following structure [ {{ "name" : name of the person you selected, "reason" : reason for selecting the person }}, ... ] # Notes - The output must be a list of JSON objects ranked in the order of preference. - You can make at most 1 selection.
ALGORITHM 1 Example prompt regarding social network data.
@@
ALGORITHM 2 General Prompt used to implement $\mathcal{Q}(A_{t},i_{t},\delta)$.
@@
We employ multiple temperatures to account for the variability in response generation by LLM systems, which is also observed in classical statistical models of network formation (Jackson, 2008). Our study conducts experiments using three temperatures for all models except Claude 3.5: 0.5, 1.0, and 1.5. For Claude 3.5 the temperature range is between 0 and 1, and we run experiments with two temperatures: 0.5, and 1.0.
@@
Feature Representations for Prompts
Principle 1: Preferential Attachment. We have the following features:
"name" : 0,
"neighbors" : [5, 7, 1, 6]
Principle 2: Triadic Closure. We have the following features:
"name" : 0,
"common_neighbors" : [5, 7, 1, 6]
Principle 3: Homophily. We have the following features:
"name" : 0,
"favorite_color" : "red",
"hobby" : "hiking",
"location" : "Boston"
Principle 5: Small-World. We have the following features:
"name" : 0,
"neighbors" : [5, 7, 1, 6]
Real-World Data. We have the following features:
"name" : 0,
"status" : "student",
"major" : 10,
"second major" : 93,
"accommodation" : "house",
"high_school" : 5,
"graduation_year" : 2008
We note that the initial Facebook100 dataset included gender information as a feature. We chose not to include gender as one of the features as it has been shown that language models exhibit gender bias (Vig et al., 2020; Kotek et al., 2023; Bordia and Bowman, 2019). An example of the prompt using real-world social network data is given at Algorithm 1.
Robustness Checks
We tried the following LLM models:
  • GPT-3.5 (gpt-3.5-turbo)
  • GPT-4o Mini (gpt-4o-mini)
  • Llama 3 (llama-3-70b-instruct)
  • Claude 3.5 Sonnet (claude-3-5-sonnet-20240620).
For each of the models except Claude 3.5 we used three temperatures: 0.5, 1.0, 1.5. For the Claude 3.5 model we used temperatures 0.5 and 1.0 (since the model does not allow temperatures above 1.0). Finally, we experimented with different environmental prompts (e.g., friendship, collaboration, community) to test prompt sensitivity.
@@
The algorithm for the altered Watts-Strogatz model is described as follows:
@@
where $i_{1},\dots,i_{T}$ are the chooser nodes (i.e., the LLM agents who want to form a link), and $j_{1},\dots,j_{T}$ are the nodes which are chosen from the alternative sets $A_{1},\dots,A_{T}$. The likelihood function is convex, and we optimize it with the L-BFGS-B method (Liu and Nocedal, 1989). The standard errors of the coefficients are approximated as $\sqrt{-H^{-1}/N}$ where $H$ is the Hessian matrix of the log-likelihood at $(\hat{\theta}_{\mathrm{PA}},\hat{\theta}_{\mathrm{TC}},\hat{\theta}_{\mathrm{H}})$ and $N$ is the number of data points (cf. (Overgoor et al., 2019; Train, 2009)).
@@
Data and code are available on GitHub at the following link:
The real-world social network data have been taken from the sources of (Traud et al., 2012) and (Yuan et al., 2018).